After a Merchant has invested 4,500 at 7% he finds another opportunity to invest at 10%. How much should he invest at the latter rate, if he wishes to make a total income of 8% on his investment
Answer by ikleyn(52925)

.
After a Merchant has invested 4,500 at 7%, he finds another opportunity to invest at 10%.
How much should he invest at the latter rate, if he wishes to make a total income of 8% on his investment.
~~~~~~~~~~~~~~~~~

Let x be the amount to invest at 10%.

Then the total interest is  0.07*4500 + 0.1x.


The equation for unknown x is this

    0.07*4500 + 0.1x = 0.08*(4500+x),

according to the problem.  Then

    0.07*4500 + 0.1x = 0.08*4500 + 0.08x,

    0.1x - 0.08x = 0.08*4500 - 0.07*4500

        0.02x    =         45

            x    =         45/0.02 = 2250.


ANSWER.  A merchant should invest additional $2250 at the rate 10% to get his goal.

Solved.